OpenStudy (anonymous):

*****medals and fans*******Amy and Stephanie were playing a number game. Amy picked a number and told Stephanie that if she added 8 to the number and multiplied the sum by 4, it would equal 12. Which equation could Stephanie use to find the number, x?

OpenStudy (anonymous):

4 + (x + 8) = 12 4(x + 8) = 12 4(x) + 8 = 12 8(x + 4) = 12

OpenStudy (anonymous):

let x be the number. then what equation would be formed

OpenStudy (anonymous):

ummmm......4(x+8)=12 right

OpenStudy (anonymous):

correct